OpenStudy (anonymous):

Which of the following is true for the equation four times d equals twenty-four? The equation represents the number that gives twenty-four when multiplied by 4. To solve, you must divide both sides by twenty-four. Checking the solution involves division. Multiplication is the inverse operation.

OpenStudy (mikurout):

all options are right.

OpenStudy (anonymous):

True: The equation represents the number that gives twenty-four when multiplied by 4. True: Multiplication is the inverse operation.

OpenStudy (anonymous):

no, to solve, you must divide both sides by 4.

OpenStudy (anonymous):

since 4d=24 divide both sides by 4 d=6

OpenStudy (anonymous):

to check, you use multiplication, not division: 4d=24 4(6)=24 24=24

OpenStudy (mikurout):

@haiimawesome All the options are right.

OpenStudy (anonymous):

thx guys

OpenStudy (anonymous):

np.

OpenStudy (anonymous):

@mikurout no they aren't

OpenStudy (mikurout):

@haiimawesome Which option are you thinking is wrong?

OpenStudy (anonymous):

To solve, you must divide both sides by twenty-four. Checking the solution involves division. THOSE TWO ARE WRONG

OpenStudy (anonymous):

because you don't divide both sides by 24 to solve 4d=24 you divide both sides by 4

OpenStudy (anonymous):

checking involves multiplication, not division, since to check that d=6, you multiply 6x4 to get 24

OpenStudy (mikurout):

@haiimawesome if you will divide both side with 24 ,then also it can be solved.

OpenStudy (mikurout):

after dividing you will have d/6 =1 so d=6

OpenStudy (anonymous):

no, that will give you: 4d=24 4/24d=1 1/6d=1

OpenStudy (anonymous):

then i guess you could divide both sides by 1/6 to get d=6

OpenStudy (anonymous):

but nobody does that...

OpenStudy (mikurout):

yes

OpenStudy (anonymous):

that isn't the proper way

OpenStudy (mikurout):

math doesnot depend on anybody.

OpenStudy (anonymous):

true, but i can guarantee that that option was false

OpenStudy (mikurout):

thats also the right way. try to understand @haiimawesome

OpenStudy (mikurout):

i am sure all the options are right. give me a valid reason so that I will think those are wrong.

OpenStudy (anonymous):

@jjthebeast89 which options were correct?

OpenStudy (mikurout):

Logically all the options are right. Are you agree with this or not? @haiimawesome

OpenStudy (anonymous):

logically i guess...wait but how about Checking the solution involves division.

OpenStudy (mikurout):

4*d =24 so d = 6 so 24/d =4

OpenStudy (anonymous):

okay. then logically they are all correct

OpenStudy (mikurout):

yes just want to say this thing.

OpenStudy (mikurout):

@haiimawesome It is not important , how you do a problem? but the important thing is what are the possible ways for doing the same problem.

OpenStudy (anonymous):

i understand your point.